A monopoly has

Price= 1000-10Q

Total Revenue= 1000-10Q2

Marginal Cost= 100+ 10Q

Q= Quantity

Find quantity and deadweight loss of monopoly


1
Expert's answer
2021-03-04T15:07:50-0500

A monopoly produces quantity at which marginal cost is equal to marginal revenue.

Thus, "MR=MC"



From the attached image, the MR curve is twice more steep than the demand curve. Also "MR=\u2206TR\/\u2206Q"

therefore:

"MR= 1000-20Q"

Equating MR and MC gives,

"1000-20Q=100+10Q"

From this equation,

"900=30Q"

Q=Quantity = 30units of output.

Deadweight loss of the monopoly(DWLM)

"=\frac {1}{2} \times (P_{0}-P_{1}) \times (Q_{1}-Q_{0})"

"P_{0}" =price at equilibrium="1000-(10\times 30)"

"P_{0}" = 700 per unit

"P_{1}" = MC="100+(10\u00d730)" = 400 per unit

"Q_{1}" = quantity on demand= 60 units

"Q_{0}" = Quantity supplied = 30units

hence,

DWLM = "\frac{1}{2} ×(700-400)×(60-30)"

=4,500